OpenStudy (anonymous):

Triangle ABC is congruent to triangle DEF. In triangle ABC, side AB measures 9, side BC measures 3x+1, and side CA measures 7. In triangle DEF, side DE measures 9, side EF measures 2x+2, and side FD measures 7. What equation would help you to solve for the side length of BC and EF? Explain your reasoning using complete sentences.

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

hihihihi, my net is good enough before I am here and you got it, right?

OpenStudy (anonymous):

I'm still struggling to understand how it was x=4 instead of 1 @Loser66

OpenStudy (loser66):

hahaha OK, you solve for x and get x =1, right? but BC = 3x+1, not x so that BC = 3*1+1= 4 and so as EF, EF = 2x+2 =2*1+2 =4

OpenStudy (anonymous):

Ahhh ok @Loser66